OpenStudy (anonymous):

Is the anti-derivative of 2x+9 = x^2?

OpenStudy (unklerhaukus):

\[\int 2x+9\text dx=x^2+9x+c\]

OpenStudy (lgbasallote):

\[\large \int 2x +9 = \int x^2?\] lol

OpenStudy (anonymous):

yes, @lgbasallote , that's my question. I wasn't sure if I'm supposed to include the 9 in my answer

OpenStudy (anonymous):

huh? now you lost me. lol

OpenStudy (lgbasallote):

\[\LARGE x^2 + 9x + c = \frac{x^3}{3} + c\] i dont get this lol...is this differential equations?

OpenStudy (anonymous):

yea i think so.. this is listed under the "Fundamental theorem of calculus"

OpenStudy (lgbasallote):

isnt the FTC \[\large \int_a^b f ' (x) dx = F(b) - F(a)\]